uniform meterstick is balanced at its midpoint with a single support. A 65 N weight is suspended at the 39 cm mark. At
what point must a 40 N weight be hung to balance the system?
1
Expert's answer
2019-06-18T08:09:30-0400

Using the law of the lever [https://en.wikipedia.org/wiki/Lever#Law_of_the_lever]

"\frac{F_b}{F_a}=\frac{a}{b},"


we get

"\frac{40}{65}=\frac{39}{b},"

from where


"b = 63.375."

Answer: 63.375 cm.